OpenStudy (anonymous):

Simplify. 2 + 4 • (6 – 3)

OpenStudy (anonymous):

2+4*(6-3) = 2+ 4*3 = 2+ 12 = 14

OpenStudy (anonymous):

18

OpenStudy (anonymous):

@alexandra 12

OpenStudy (anonymous):

@Nick123cross1 how did you get 18?

OpenStudy (anonymous):

it's 12, sorry forgot add